OpenStudy (anonymous):

A rock group plans three concert tours over a period of 32 weeks. The tour in Britain will be 2 weeks longer than the tour in France and the tour in Germany will be 6 weeks shorter than the tour in France. How many weeks will they be in France?

OpenStudy (unklerhaukus):

12

OpenStudy (unklerhaukus):

b+g+f=32 b=f+2 g=f-6

OpenStudy (unklerhaukus):

f+2 + f-6 + f = 32 3f-4=32 3f=36 f=12